# Bumblebee gravity: spherically-symmetric solutions away from the potential minimum

### v1

@@ IV.4 Some specialized exact solutions
Some special solutions can be countenanced for the equations in (49). They stem from assuming the condition $(AB)^{\prime}=0$. This condition implies $AB=\alpha$ where $\alpha$ is a constant. By multiplying (48) by $AB$, we see that
$-2\kappa V^{\prime}B^{2}B_{0}^{2}r=0$ (55)
@@
where the $c_{n}$’s are now distinct coefficients. Because $V^{\prime}=0$, we must have that the argument of $V$ is constant, and thus we can let $B_{0}^{2}/A=-\beta$, where $\beta$ is a nonnegative constant, taking the value of one of the minima of the potential. For example, for the hypergeometric case, from equation (54), $\beta=3$, and $\bar{V}$ is given by (41).
@@
$\bigg{(}-\frac{\kappa\alpha\beta{\bar{V}}}{3}\bigg{)}r^{2}+\bigg{(}c_{1}^{2}+\alpha\beta\bigg{)}+\bigg{(}c_{3}\beta-2c_{1}c_{2}\bigg{)}\frac{1}{r}+\bigg{(}c_{2}^{2}-\frac{\kappa\beta c_{2}^{2}}{2}\bigg{)}\frac{1}{r^{2}}=0.$ (60)
@@
$B_{0}$ =±$=\pm\sqrt{-\alpha\beta},$ (62)
@@
This solution is Minkowksi spacetime. If $\beta=2/\kappa$, we see that $c_{3}=\kappa c_{2}\sqrt{-\alpha\beta}$, and we arrive at the solution
$B_{0}$ =±$=\pm\left(\sqrt{-\alpha\beta}-\frac{c_{2}}{r}\right),$ (63)
@@
which is of the RN form (we add a $\pm$ to the $c_{2}$ term in $B_{0}$ without loss of generality and to emphasize that $B_{0}^{2}=-\beta A$). To match with the standard form $A=-1+2GM/r-GQ^{2}/(4\pi r^{2})$, $\alpha=-1$, $c_{2}=\pm Q/(4\pi)$, and charge and mass are related by $Q=\mp M/\sqrt{\beta}$. Since a nonzero $Q$ arises through the potential, this shows that a RN solution could arise in the absence of actual electric charge, as in Ref. [35], and others. Note that the constraint ${\bar{V}}=0$ can severely limit the possible form of the potential $V(X)$.
@@ V Coupled Gravity-Vector Solutions, Part II
The transformation between EF and Schwarzschild coordinates ($t,r,\theta,\phi$) is accomplished by $dt=dv-(M/N)dr$, so that with (45), $B=M^{2}/N$, and $A=-N$. A primary utility of these coordinates for our purposes is that at a horizon the functions $N$, and $M$ do not diverge, rather $N(r_{h})=0$ at a horizon $r_{h}$. Also in these coordinates the vector field takes the form $B_{\mu}=(B_{v},B_{r},0,0)$.
@@
$B_{v}^{\prime\prime}+\frac{2}{r}B_{v}^{\prime}$ $=2V^{\prime}\frac{M}{N}B_{v}+\kappa V^{\prime}\frac{M^{2}}{N^{2}}B_{v}^{2}B^{\prime}_{v}r,$ (65)
@@
$f^{\prime\prime}+\frac{2}{x}f^{\prime}$ $=2{\tilde{V}}^{\prime}\frac{M}{N}f+k{\tilde{V}}^{\prime}\frac{M^{2}}{N^{2}}f^{2}f^{\prime}x,$ (66)
@@
The system of equations in (66) are second order non-autonomous (an explicit $x$ dependence) ordinary differential equations. For $x>0$ and $N\neq 0$, the solutions are unique once initial values for $f$, $f^{\prime}$, $M$, and $N$ are given at some $x_{0}\neq 0$ [116]. The system (66) can be made first order by using $z=f^{\prime}$ to yield a first order equation. It is still technically non-autonomous so many of the ODE techniques for characterizing solutions based on integral curves for example will not apply (at least not in a straightforward manner).
The arguments for the global behavior of the function $f$ made in the previous section III.2.3 are no longer straightforward due to the extra term proportional to the gravity coupling $k$ on the right-hand side of the $f$ equation in (66). This term is proportional to $f^{\prime}$ and hence it is difficult makes make arguments about its sign. However, one can discern the behavior of $M$ more easily than the $N$ or $f$. Assuming $M>0$, all the terms on the right-hand side are positive save ${\tilde{V}}^{\prime}$. Thus the equation for $M^{\prime}$ tells us that
@@ V.1.1 Horizon
We consider first the assumption of a regular horizon. In EF coordinates this means $N(1)=0$ and $N(x),M(x),f(x)$ are all analytical functions in the neighborhood of $x=1$ ($r=r_{h}$). We form series expansions for these functions in terms of $\delta x=x-1$:
@@
where $m$ is the desired truncation. By construction, this expansion must have $a_{N}(0)=0$. The expansions (68) are inserted into the ODEs (66), and the coefficients are fixed order by order in $\delta x^{j}$ up to the desired term truncation. Running through this calculation we find computation time is saved if we carry out this calculation to obtain an expansion to order $\delta x^{5}$ or $\delta x^{6}$, which, due to the derivatives and small denominator terms, requires at least $m=8$ in (68).
@@
$N(x)$ $=(1-\tfrac{1}{2}a_{f}(1)^{2})\delta x+(ka_{f}(1)^{2}-1)\delta x^{2}+...,$
where the solutions depend only on the horizon values $a_{f}(0)$ and $a_{f}(1)$. Clearly, the type of potential ${\tilde{V}}$ chosen influences the relations among the coefficients $a_{f}(j)$, $a_{M}(j)$, and $a_{N}(j)$. For any choice ${\tilde{V}}\neq 0$ of potential, we find $a_{f}(0)=0$ and the coefficient $a_{f}(1)$ is the only free parameter left, fixed by the value of $df/dx$ at the Horizon. To capture the dependencies of the coefficients, we show them explicitly for the first 3 terms in the series in Tables 2 and 3.
@@
0 $a_{f0}$ $1$ 00 00 $1$ 00
1 $a_{f1}$ 00 $\tfrac{1}{2}\alpha_{1}$ $a_{f1}$ $2{\tilde{\mu}}^{2}a_{f1}^{2}/\alpha_{1}^{2}$ $\tfrac{1}{2}\alpha_{1}$
2 $-a_{f1}$ 00 $-\alpha_{2}$ $a_{f1}\left(\frac{2{\tilde{\mu}}^{2}}{\alpha_{1}^{2}}-1\right)$ $\frac{{\tilde{\mu}}^{2}a_{f1}^{2}[4-8a_{f1}^{2}+3a_{f1}^{4}+4{\tilde{\mu}}^{2}(1+a_{f1}^{2})]}{\alpha_{1}^{4}}$ $\frac{[(6+{\tilde{\mu}}^{2})a_{f1}^{2}-2a_{f1}^{4}-4]}{2\alpha_{1}}$
Table 2: First 3 terms in the horizon analytic expansion (68) for the RN and massive potential case. We take $k=1$, $a_{M}(0)=1$. Also, $\alpha_{1}=2-a_{f1}^{2}$, $\alpha_{2}=1-a_{f1}^{2}$ and we abbreviate $a_{f1}=a_{f}(1)$.
@@ 0 $a_{f0}$ $1$ 00 00 $1$ 00
1 $a_{f1}$ $\frac{4{\tilde{\lambda}}a_{f1}^{2}}{({\tilde{\lambda}}-\alpha_{1})^{2}}$ $\tfrac{1}{2}(\alpha_{1}-{\tilde{\lambda}})$
2 $-a_{f1}\frac{[3{\tilde{\lambda}}^{2}-2{\tilde{\lambda}}(2+\alpha_{1})+\alpha_{1}^{2}]}{({\tilde{\lambda}}-\alpha_{1})^{2}}$ $\frac{2{\tilde{\lambda}}a_{f1}^{2}[4-5{\tilde{\lambda}}^{2}-12a_{f1}^{2}+5a_{f1}^{4}+4{\tilde{\lambda}}(2+3a_{f1}^{2})]}{({\tilde{\lambda}}-\alpha_{1})^{4}}$ $\frac{(2-{\tilde{\lambda}}-3a_{f1}^{2}+a_{f1}^{4})}{{\tilde{\lambda}}-\alpha_{1}}$
$M(3,2,X/\Lambda^{2})$
@@
0 00 $1$ 00
1 $a_{f1}$ $6e^{3}{\tilde{g}}a_{f1}^{2}/\beta_{1}^{2}$ $-\beta_{1}/(2e^{3})$
2 $-a_{f1}\left(1+\frac{3e^{3}{\tilde{g}}\beta_{2}}{\beta_{1}^{2}}\right)$ $\frac{e^{6}{\tilde{g}}a_{f1}^{2}[-3\beta_{2}({\tilde{g}}+e^{3}(2+8{\tilde{g}}))+2e^{3}(7{\tilde{g}}+4e^{3}(8{\tilde{g}}-5))a_{f1}^{2}+17e^{6}a_{f1}^{4}]}{\beta_{1}^{4}}$ $\frac{-2{\tilde{g}}\alpha_{2}+e^{3}(4-4{\tilde{g}}+({\tilde{g}}-6)a_{f1}^{2}+2a_{f1}^{4})}{2\beta_{1}}$
@@ 1 $a_{f1}$ $6e^{3}{\tilde{g}}a_{f1}^{2}/\beta_{1}^{2}$ $-\beta_{1}/(2e^{3})$
0 00 $1$ 00
1 $a_{f1}$ $72e^{4}{\tilde{g}}a_{f1}^{2}/\beta_{3}^{2}$ $-\beta_{3}/6e^{2}$
2 $-a_{f1}\left(1+\frac{24e^{2}g\beta_{4}}{\beta_{3}^{2}}\right)$ $\frac{36e^{4}{\tilde{g}}a_{f1}^{2}[57e^{4}a_{f1}^{4}-4\beta_{4}({\tilde{g}}+3e^{2}(1+5{\tilde{g}}))+4e^{2}(8{\tilde{g}}+e^{2}(60{\tilde{g}}-33))a_{f1}^{2}]}{\beta_{3}^{4}}$ $\frac{-2{\tilde{g}}\alpha_{2}+3e^{2}(2-2{\tilde{g}}-3a_{f1}^{2}+a_{f1}^{4})}{\beta_{3}}$
Table 3: The first 3 terms in the horizon analytic expansion (68) for the quadratic and 2 sample hypergeometric potentials. We take $k=1$, $a_{M}(0)=1$. Also, we use abbreviations: $\beta_{1}={\tilde{g}}+e^{3}(2{\tilde{g}}-\alpha_{1})$, $\beta_{2}={\tilde{g}}+2e^{3}({\tilde{g}}-1)$, $\beta_{3}=2{\tilde{g}}+3e^{2}(2{\tilde{g}}-\alpha_{1})$, and $\beta_{4}={\tilde{g}}+3e^{2}({\tilde{g}}-1)$.
From the table one can see some general features. As a guide, we focus on the horizon values that yield an RN solution that asymptotes to a constant value. For this behavior, from the RN entry in Table 2, we would need $\alpha_{1}>0$, so that $N$ is increasing from 00 and we need to ensure the second derivative is negative, or $-\alpha_{2}<0$. These two criteria can be achieved if $a_{f1}^{2}<1$. Given these criteria, we examine the behavior implied by the four potential cases in the Tables 2 and 3. For the massive potential, the behavior does depend on the size of ${\tilde{\mu}}$. Looking at the behavior of $N$, we see that $dN/dx>0$ if $\alpha_{1}>0$. If $a_{f1}=0$, the solution appears to reduce to the Schwarschild behavior with the vector field vanishing (this is consistent with known results). If $a_{f1}^{2}<1$ (but nonzero), then one can show that $d^{2}N/dx^{2}>0$ for all nonzero ${\tilde{\mu}}$ which can yield non-asymptotically flat behavior. The behavior of $f(x)$ is determined by the sign of $a_{f1}$, while $M$ will always be initially increasing with concavity of varying sign depending on ${\tilde{\mu}}$.
@@ 1 $a_{f1}$ $72e^{4}{\tilde{g}}a_{f1}^{2}/\beta_{3}^{2}$ $-\beta_{3}/6e^{2}$
For the hypergeometric potentials it is more challenging to discern which choices of the coupling ${\tilde{g}}$ and the value $a_{f1}$ will yield $dN/dx>0$ and $N^{\prime\prime}<0$ as in the RN case. Some studies of the dependence of $N^{\prime}$ and $N^{\prime\prime}$ on these parameters for the $M(3,2,X)$ case and the $M(4,2,X)$, reveals that for a range of values of described by approximately ${\tilde{g}}^{2}+a_{f1}^{2}<1$ (with $a_{f1}>0$ and ${\tilde{g}}>0$), then $N^{\prime}>0$ and $N^{\prime\prime}<0$ at the horizon. So it is possible for a solution to initiate behavior similar to the RN case. However, the solution for $M$ in this case has $M^{\prime}>0$ and $M^{\prime\prime}>0$, indicating a possible divergence.
These remarks only apply for $x$ near the Horizon ($\delta x<<1$). Ultimately, the “exact” solution will show whether solutions mimic the RN behavior or not as $x$ increases. The next step is to use the near-horizon behavior to generate initial conditions for precise numerical solutions. This is carried out in Mathematica [119]. We also carry out the horizon expansion up to 5th or 6th order in $\delta x$. The higher coefficients $a_{f}(j)$, $a_{M}(j)$, and $a_{N}(j)$ for $j>2$ are much too lengthy to record here but are stored in Mathematica files. For a numerical solver, the differential equations (66) are singular at $N=0$ ($x=1$) so we step away from this point slightly ($x=1.05$) and find initial conditions from the horizon expansion.
In the plots to follow, we show both the horizon expansion, and the numerical results together for several choices of potential, and sample initial values at the horizon. No approximation is made for the numerical solutions; the full equations are used. The numerical scheme, employing “NDSolveValue” in Mathematica, uses a combination of varying order Rung-Kutta techniques designed to carefully approach stiff or singular points. One can see where the horizon expansion breaks down in these plots. A common feature we discern is that for various choices of potential, the full numerical code appears to carve out quite a different path from the initial trends indicated by the horizon expansion.
@@
(b) The quadratic potential case. The series solution is truncated to $\delta x^{5}$. The initial value is $f^{\prime}(1)=-1/3$ and ${\tilde{\lambda}}=1/5$ at the horizon.
@@
(c) The hypergoemetric potential $M(3,2,X)$ case. The series solution is truncated to $\delta x^{6}$. The initial value used is $f^{\prime}(1)=-1/10$, while and ${\tilde{g}}=1/10$.
@@
(d) The hypergeometric potential $M(4,2,X)$ case. The series solution is truncated to $\delta x^{6}$. The initial value used is $f^{\prime}(1)=-1/9$, while and ${\tilde{g}}=1/8$. The peak for the $M$ curve extends to $M\approx 7.5$.
@@
Figure 6: Numerical Solutions using the Horizon expansion (68) to generate initial values near $x=1$. The analytical results with the series solution are labeled with subscript $A$. The dashed lines are the RN case while the dotted lines show the analytical series from (68). The connection with conventional Schwarzschild coordinates is $g_{tt}=A=-N$
@@
The first plot 6(a) is the Massive vector potential case. The plot shows that when there is a horizon, the fields $N$, $M$, $f$ show diverging behavior - which occurs in the numerical simulation. This implies that there is either another horizon or a physical singularity in the spacetime. The situation is resolved by checking the curvature in Section V.1.3, which shows that this solution, which starts from a horizon, leads to a physical singularity. Thus the solution is not asymptotically flat; consistent with the Bekenstein result.
@@
Consider next the Quadratic potential case in the second plot 6(b). This plot shows that when the solution starts from a horizon at $x=1$, there is an interesting pulse-like behavior in $M$ at about $r\approx 4r_{h}$, followed by a decrease to zero for $N$ and $M$. Again, we can ask the nature of the singularity as $N\rightarrow 0$. The answer provided by the curvature is that as $x\rightarrow\infty$, scalar curvature quantities appear to diverge, suggesting a singularity.

### v2

@@ IV.4 Some specialized exact solutions
Some special solutions can be countenanced for the equations in (49). They stem from assuming the condition $(AB)^{\prime}=0$. This condition implies $AB=\alpha$ where $\alpha$ is a constant. By multiplying equation (48) by $AB$, we see that
$2\kappa V^{\prime}B^{2}B_{0}^{2}r=0.$ (55)
@@
where the $c_{n}$’s are now distinct coefficients. Because $V^{\prime}=0$, we must have that the argument of $V$ is constant, and thus we can let $B_{0}^{2}/A=-\beta$, where $\beta$ is a non-negative constant, taking the value of one of the minima of the potential. For example, for the hypergeometric case, from equation (54), $\beta=3\Lambda^{2}$, and $\bar{V}$ is given by (41).
@@
$\bigg{(}-\frac{\kappa\alpha\beta{\bar{V}}}{3}\bigg{)}r^{2}+\bigg{(}c_{1}^{2}+\alpha\beta\bigg{)}+\bigg{(}c_{3}\beta-2c_{1}c_{2}\bigg{)}\frac{1}{r}$ (60)
$\phantom{sp}+\bigg{(}c_{2}^{2}-\frac{\kappa\beta c_{2}^{2}}{2}\bigg{)}\frac{1}{r^{2}}=0.$
@@
$B_{0}$ $=\pm\sqrt{-\alpha\beta},$ (62)
@@
This solution is Minkowski spacetime. If $\beta=2/\kappa$, we see that $c_{3}=\kappa c_{2}\sqrt{-\alpha\beta}$, and we arrive at the solution
$B_{0}$ $=\pm\left(\sqrt{-\alpha\beta}-\frac{c_{2}}{r}\right),$ (63)
@@
which is of the RN form (we add a $\pm$ to the $c_{2}$ term in $B_{0}$ without loss of generality and to emphasize that $B_{0}^{2}=-\beta A$). To match with the standard form $A=-1+2GM/r-GQ^{2}/(4\pi r^{2})$, $\alpha=-1$, $c_{2}=\pm Q/(4\pi)$, and charge and mass are related by $Q=\pm M/\sqrt{\beta}$. Since a nonzero $Q$ arises through the potential, this shows that a RN solution could arise in the absence of actual electric charge, as in Ref. [35], and others. Note that the constraint ${\bar{V}}=0$ can severely limit the possible form of the potential $V(X)$.
@@ V Coupled Gravity-Vector Solutions, Part II
The transformation between EF and Schwarzschild coordinates ($t,r,\theta,\phi$) is accomplished by dt=dv(M/N)drdt=dv-(M/N)dr, so that with (45), $B=M^{2}/N$, and A=NA=-N. A primary utility of these coordinates for our purposes is that at a horizon the functions $N$, and $M$ do not diverge, rather $N(r_{h})=0$ at a horizon $r_{h}$. Also in these coordinates the vector field takes the form $B_{\mu}=(B_{v},B_{r},0,0)$.
@@
$B_{v}^{\prime\prime}+\frac{2}{r}B_{v}^{\prime}$ $=2V^{\prime}\frac{M^{2}}{N}B_{v}+\kappa V^{\prime}\frac{M^{2}}{N^{2}}B_{v}^{2}B^{\prime}_{v}r,$ (65)
@@
$f^{\prime\prime}+\frac{2}{x}f^{\prime}$ $=2{\tilde{V}}^{\prime}\frac{M^{2}}{N}f+k{\tilde{V}}^{\prime}\frac{M^{2}}{N^{2}}f^{2}f^{\prime}x,$ (66)
@@
The system of equations in (66) are second order nonautonomous (an explicit $x$ dependence) ordinary differential equations. For $x>0$ and $N\neq 0$, the solutions are unique once initial values for $f$, $f^{\prime}$, $M$, and $N$ are given at some $x_{0}\neq 0$ [116]. The system (66) can be made first order by using $z=f^{\prime}$ to yield a first order equation. It is still technically nonautonomous so many of the ODE techniques for characterizing solutions based on integral curves for example will not apply (at least not in a straightforward manner).
The arguments for the global behavior of the function $f$ made in the previous section III.2.3 are no longer straightforward due to the extra term proportional to the gravity coupling $k$ on the right-hand side of the $f$ equation in (66). This term is proportional to $f^{\prime}$ and hence it is difficult to make arguments about its sign. However, one can discern the behavior of $M$ more easily than the $N$ or $f$. Assuming $M>0$, all the terms on the right-hand side are positive save $\tilde{V}$$^{\prime}$. Thus the equation for $M^{\prime}$ tells us that
@@ V.1.1 Horizon
We consider first the assumption of a regular horizon. In EF coordinates this means $N(1)=0$ and $N(x),M(x),f(x)$ are all analytic functions in the neighborhood of $x=1$ ($r=r_{h}$). We form series expansions for these functions in terms of $\delta x=x-1$:
@@
where $m$ is the desired truncation. By construction, this expansion must have $a_{N}(0)=0$. The expansions (68) are inserted into the ODEs (66), and the coefficients are fixed order by order in $\delta x^{j}$ up to the desired term truncation. Running through this calculation we find computation time is saved if we carry out this calculation to obtain an expansion to order $\delta x^{5}$ or $\delta x^{6}$, which, due to the derivatives and small denominator terms, requires at least $m=8$ in equation (68).
@@
$N(x)$ $=(1-\tfrac{1}{2}a_{f}(1)^{2})\delta x+(ka_{f}(1)^{2}-1)\delta x^{2}+$.,
where the solutions depend only on the horizon values $a_{f}(0)$ and $a_{f}(1)$. Clearly, the type of potential $\tilde{V}$} chosen influences the relations among the coefficients $a_{f}(j)$, $a_{M}(j)$, and $a_{N}(j)$. For any choice ${\tilde{V}}\neq 0$ of potential, we find $a_{f}(0)=0$ and the coefficient $a_{f}(1)$ is the only free parameter left, fixed by the value of $df/dx$ at the horizon. To capture the dependencies of the coefficients, we show them explicitly for the first three terms in the series in Tables 2 and 3.
Table 2: First three terms in the horizon analytic expansion (68) for the RN and massive potential cases. We take $k=1$, $a_{M}(0)=1$. Also, $\alpha_{1}=2-a_{f1}^{2}$, $\alpha_{2}=1-a_{f1}^{2}$ and we abbreviate $a_{f1}=a_{f}(1)$.
@@
0 $a_{f0}$ 11 0 0 11 0
1 $a_{f1}$ 0 $\tfrac{1}{2}\alpha_{1}$ $a_{f1}$ $2{\tilde{\mu}}^{2}a_{f1}^{2}/\alpha_{1}^{2}$ $\tfrac{1}{2}\alpha_{1}$
2 $a_{f1}$ 0 $\alpha_{2}$ $a_{f1}\left(\frac{2{\tilde{\mu}}^{2}}{\alpha_{1}^{2}}-1\right)$ $\frac{{\tilde{\mu}}^{2}a_{f1}^{2}[4-8a_{f1}^{2}+3a_{f1}^{4}+4{\tilde{\mu}}^{2}(1+a_{f1}^{2})]}{\alpha_{1}^{4}}$ $\frac{[(6+{\tilde{\mu}}^{2})a_{f1}^{2}-2a_{f1}^{4}-4]}{2\alpha_{1}}$
Table 3: The first three terms in the horizon analytic expansion (68) for the quadratic and two sample hypergeometric potentials. We take $k=1$, $a_{M}(0)=1$. Also, we use abbreviations: $\beta_{1}={\tilde{g}}+e^{3}(2{\tilde{g}}-\alpha_{1})$, $\beta_{2}={\tilde{g}}+2e^{3}({\tilde{g}}-1)$, $\beta_{3}=2{\tilde{g}}+3e^{2}(2{\tilde{g}}-\alpha_{1})$, and $\beta_{4}={\tilde{g}}+3e^{2}({\tilde{g}}-1)$.
@@ 0 $a_{f0}$ 11 0 0 11 0
[…]
2 $a_{f1}\frac{[3{\tilde{\lambda}}^{2}-2{\tilde{\lambda}}(2+\alpha_{1})+\alpha_{1}^{2}]}{({\tilde{\lambda}}-\alpha_{1})^{2}}$ $\frac{2{\tilde{\lambda}}a_{f1}^{2}[4-5{\tilde{\lambda}}^{2}-12a_{f1}^{2}+5a_{f1}^{4}+4{\tilde{\lambda}}(2+3a_{f1}^{2})]}{({\tilde{\lambda}}-\alpha_{1})^{4}}$ $\frac{(2-{\tilde{\lambda}}-3a_{f1}^{2}+a_{f1}^{4})}{{\tilde{\lambda}}-\alpha_{1}}$
$M(3,2,z)$
@@
0 0 11 0
1 $a_{f1}$ $6e^{6}{\tilde{g}}a_{f1}^{2}/\beta_{1}^{2}$ $\beta_{1}/(2e^{3})$
2 $a_{f1}\left(1+\frac{3e^{3}{\tilde{g}}\beta_{2}}{\beta_{1}^{2}}\right)$ $\frac{e^{6}{\tilde{g}}a_{f1}^{2}[-3\beta_{2}({\tilde{g}}+e^{3}(2+8{\tilde{g}}))+2e^{3}(7{\tilde{g}}+4e^{3}(8{\tilde{g}}-5))a_{f1}^{2}+17e^{6}a_{f1}^{4}]}{\beta_{1}^{4}}$ $\frac{-2{\tilde{g}}\alpha_{2}+e^{3}(4-4{\tilde{g}}+({\tilde{g}}-6)a_{f1}^{2}+2a_{f1}^{4})}{2\beta_{1}}$
@@ 1 $a_{f1}$ $6e^{6}{\tilde{g}}a_{f1}^{2}/\beta_{1}^{2}$ $\beta_{1}/(2e^{3})$
0 0 11 0
1 $a_{f1}$ $72e^{4}{\tilde{g}}a_{f1}^{2}/\beta_{3}^{2}$ $\beta_{3}/6e^{2}$
2 $a_{f1}\left(1+\frac{24e^{2}g\beta_{4}}{\beta_{3}^{2}}\right)$ $\frac{36e^{4}{\tilde{g}}a_{f1}^{2}[57e^{4}a_{f1}^{4}-4\beta_{4}({\tilde{g}}+3e^{2}(1+5{\tilde{g}}))+4e^{2}(8{\tilde{g}}+e^{2}(60{\tilde{g}}-33))a_{f1}^{2}]}{\beta_{3}^{4}}$ $\frac{-2{\tilde{g}}\alpha_{2}+3e^{2}(2-2{\tilde{g}}-3a_{f1}^{2}+a_{f1}^{4})}{\beta_{3}}$
From the table one can see some general features. As a guide, we focus on the horizon values that yield an RN solution that asymptotes to a constant value. For this behavior, from the RN entry in Table 2, we would need $\alpha_{1}>0$, so that $N$ is increasing from 0 and we need to ensure the second derivative is negative, or $\alpha_{2}<0$. These two criteria can be achieved if $a_{f1}^{2}<1$. Given these criteria, we examine the behavior implied by the four potential cases in Tables 2 and 3. For the massive potential, the behavior does depend on the size of $\tilde{\mu}$}. Looking at the behavior of $N$, we see that $dN/dx>0$ if $\alpha_{1}>0$. If $a_{f1}=0$, the solution appears to reduce to the Schwarzschild behavior with the vector field vanishing (this is consistent with known results). If $a_{f1}^{2}<1$ (but nonzero), then one can show that $d^{2}N/dx^{2}>0$ for all nonzero $\tilde{\mu}$} which can yield nonasymptotically flat behavior. The behavior of $f(x)$ is determined by the sign of $a_{f1}$, while $M$ will always be initially increasing with concavity of varying sign depending on $\tilde{\mu}$}.
@@ 1 $a_{f1}$ $72e^{4}{\tilde{g}}a_{f1}^{2}/\beta_{3}^{2}$ $\beta_{3}/6e^{2}$
For the hypergeometric potentials it is more challenging to discern which choices of the coupling $\tilde{g}$} and the value $a_{f1}$ will yield $dN/dx>0$ and $N^{\prime\prime}<0$ as in the RN case. Some studies of the dependence of $N^{\prime}$ and $N^{\prime\prime}$ on these parameters for the $M(3,2,z)$ case and the $M(4,2,z)$, reveal that for a range of values described approximately by ${\tilde{g}}^{2}+a_{f1}^{2}<1$ (with $a_{f1}>0$ and ${\tilde{g}}>0$), then $N^{\prime}>0$ and $N^{\prime\prime}<0$ at the horizon. So it is possible for a solution to initiate behavior similar to the RN case. However, the solution for $M$ in this case has $M^{\prime}>0$ and $M^{\prime\prime}>0$, indicating a possible divergence.
These remarks only apply for $x$ near the horizon ($\delta x<<1$). Ultimately, the “exact” solution will show whether solutions mimic the RN behavior or not as $x$ increases. The next step is to use the near-horizon behavior to generate initial conditions for precise numerical solutions. This is carried out in Mathematica [119]. We also carry out the horizon expansion up to fifth or six order in $\delta x$. The higher coefficients $a_{f}(j)$, $a_{M}(j)$, and $a_{N}(j)$ for $j>2$ are much too lengthy to record here but are stored in Mathematica files. For a numerical solver, the differential equations (66) are singular at $N=0$ ($x=1$) so we step away from this point slightly ($x=1.05$) and use initial conditions from the horizon expansion at that point.
In the plots to follow, we show both the horizon expansion, and the numerical results together for several choices of potential, and sample initial values at the horizon. No approximation is made for the numerical solutions; the full equations are used. The numerical scheme, employing “NDSolveValue” in Mathematica, uses a combination of varying order Runge-Kutta techniques designed to carefully approach stiff or singular points. One can see where the horizon expansion breaks down in these plots. A common feature we discern is that for various choices of potential, the full numerical code appears to carve out quite a different path from the initial trends indicated by the horizon expansion.
@@
(b) The quadratic potential case. The series solution is truncated to $\delta x^{5}$. The initial value is $f^{\prime}(1)=-1/5$ and ${\tilde{\lambda}}=1/5$ at the horizon.
@@
(c) The hypergoemetric potential $M(3,2,z)$ case. The series solution is truncated to $\delta x^{6}$. The initial value used is $f^{\prime}(1)=-1/10$, while and ${\tilde{g}}=1/10$.
@@
(d) The hypergeometric potential $M(4,2,z)$ case. The series solution is truncated to $\delta x^{5}$. The initial value used is $f^{\prime}(1)=-1/9$, while and ${\tilde{g}}=1/8$. The peak for the $M$ curve extends to $M\approx 7.5$.
@@
Figure 6: Numerical solutions using the horizon expansion (68) to generate initial values near $x=1$. The analytical results with the series solution are labeled with subscript AA. The dashed lines are the RN case while the dotted lines show the analytical series from equation (68). The connection with conventional Schwarzschild coordinates is $g_{tt}=A=-N$ and $z=X/\Lambda^{2}$.
@@
The first plot Fig. 6a, is the massive vector potential case. The plot shows that when there is a horizon, the fields $N$, $M$, $f$ show diverging behavior - which occurs in the numerical simulation. This implies that there is either another horizon or a physical singularity in the spacetime. The situation is resolved by checking the curvature in Section V.1.3, which shows that this solution, which starts from a horizon, leads to a physical singularity. Thus the solution is not asymptotically flat; consistent with the Bekenstein result.
@@
Consider next the quadratic potential case in the second plot Fig. 6b. This plot shows that when the solution starts from a horizon at $x=1$, there is an interesting pulse like behavior in $M$ at about $r\approx 4r_{h}$, followed by a decrease to zero for $N$ and $M$. Again, we can ask the nature of the singularity as $N\rightarrow 0$. The answer provided by the curvature is that as $x\rightarrow\infty$, scalar curvature quantities appear to diverge, suggesting a singularity.